C= 85+0.75yd I = 30 G= 200 T = 20+0.2y

Yd = Y-T

Find the equilibrium level of income.

Calculate the equilibrium level of saving and consumption.

Find the value of APC and APS .


Y= C+I+G+NX

NX= X-M

C= 50+0.8Y I = 50 G = 100 X = 250 M = 0.2y + 100

Find the equilibrium level of National Income.

Find the Consumption level , Import or saving


1
Expert's answer
2020-12-15T11:24:42-0500

1) The equilibrium level of income is:

"Y = C + I + G = 85+0.75(Y - 20 - 0.2Y) + 30 + 200,"

0.4Y = 300,

Y = 750,

Consumption is:

"C = 70 + 0.6\u00d7750 = 520,"

S = Y - C = 750 - 520 = 230.

"APC = 520\/750 = 0.693,"

APS = 1 - APC = 0.307.

2) "Y = C + I + G + NX = 50 + 0.8Y + 50 + 100 + (250 - 0.2Y - 100),"

0.4Y = 350,

Y = 875.

"C = 50 + 0.8\u00d7875 = 750,"

S = 875 - 750 = 125.